# Towards Open-World Retrieval-Augmented Generation on Knowledge Graph: A Multi-Agent Collaboration Framework

### v1

@@ 3.3 Anchor Entity Identifying
where $\mathcal{N}(e)$ denotes the set of top-$k$ relevant relations according to the Eq.(1). By scoring and ranking all candidates in the set $\tilde{\mathcal{E}}$, the predictor agent selects the top-$m$ entities with the higher relevance scores (e.g., France and Nijmegen) as the final anchor candidate set, which will be inputted to the following retriever agents for parallel searching.
3.4 Rough Pruning and Fine Filtering
Given the top-$m$ candidate anchor entities obtained from predictor agent, we assign each entity to an independent retriever agent. Thus there exist $m$ retrieval agents operating in parallel, each conducting iterative retrieval and reasoning based on its assigned anchor entity. At each iteration, these agents will retrieve the one-hop neighborhood of the given entity, where the neighborhood includes the neighboring relations and entities. Then a rough pruning and fine filtering method is applied to select more meaningful neighborhood for the question. Finally, the neighboring entities will be considered as the initial ones for the next retrieval iteration if necessary.
Neighboring Relations Pruning:
At the $t$-th retrieval iteration, each retriever agent explores candidate relations from its current given entities $e_{i}^{t}$. This agent retrieves all one-hop incoming and outgoing relations of the given entity from the knowledge graph $\mathcal{G}$. To reduce noise, the agent will filter out trivial or schema-level relations (e.g., type.object.name, common.*, sameAs), as well as previously visited relations. The remaining relations are scored by the LLM according to their relevance score with the given question. Specifically, each agent constructs a prompt containing the question, current anchor entity, and candidate relations, applying the LLM to assign the normalized scores. The prompt and in-context examples can be found in the Appendix. With the neighboring relations pruning, the top-$b$ relations are selected.
Neighboring Entities Pruning:
Given the selected relations, each agent retrieves the neighboring entities linked via these relations. Considering that each relation will have many neighboring entities, to bound the length of LLM input, we limit the number of entities per relation. Specifically, each candidate entity is ranked by the LLM with the prompt including the question, current anchor entity and the selected relation. The prompt and in-context examples can be found in the Appendix. Then top-$b$ entities are obtained for each relation, forming the knowledge triple set $\mathcal{T}(e_{i}^{t})$.
Candidate Triples Filtering:
To further identify the question-related facts implied in the knowledge graph, each agent applies a fine filtering method on the candidate triples $\mathcal{T}(e_{i}^{t})$, improving the performance in the real-world question answering. Specifically, the retrieval agent first integrates all candidate triples in $\mathcal{T}(e_{i}^{t})$ into a structured list and prompts a LLM to select the related triples according to the given question. The prompt and in-context examples can be found in the Appendix. Then the filtered triples are adopted as high-quality external knowledge which will be inputted to the following supervisor agent for answer generation. Furthermore, if the filtered triple set is empty, it means that the current exploration path can no longer yield related information for the given question. In that case, this agent will terminate its subsequent retrieval and provide no information for the supervisor at the current iteration. This “early-stopping” mechanism ensures the agent can adaptively waive unpromising knowledge paths, reducing the computational overhead for efficient retrieval-augmented generation.
@@ 3.4 Rough Pruning and Fine Filtering
Table 1: The results of different methods on various Datasets, using Qwen-Plus and GPT-4o-mini as the LLM. The best results are highlighted in bold. The suboptimal results are underlined.
3.5 Answer Generating
At the end of each iteration, all active retrieval agents input their resulting paths to a supervisor agent. The supervisor aggregates these paths and prompts the LLM to assess whether the current information is sufficient to answer the question. If so, it stops the reasoning process and returns the final answer. If not, only agents with valid paths are permitted to perform the next retrieval round. The prompt and in-context examples can be found in the Appendix. To avoid excessive computation overhead, the supervisor adopts the early termination if (1) no agents remain active, or (2) the maximum reasoning depth $L$ is reached. In such cases, the supervisor invokes LLM to directly answer the question by applying CoT reasoning with its internal knowledge.
4 Experiments
4.1 Dataset and Evaluation
@@
In the experiments, we adopt four benchmark datasets to evaluate the performance of our framework, including three multi-hop Knowledge Graph Question Answering (Multi-Hop KGQA) datasets: WebQuestionSP (WebQSP) (Yih et al. 2016), GrailQA (Gu et al. 2021) and Complex WebQuestions (CWQ) (Talmor and Berant 2018), and one for Open-Domain Question Answering: WebQuestions (Berant et al. 2013). Freebase (Bollacker et al. 2008) is utilized as the external Knowledge Graph, which contains the complete knowledge facts that are necessary to answer questions in the above datasets. In addition, we apply the exact match (i.e., Hit@1) and accuracy (i.e., acc) as evaluation metrics for our method which are widely used in the field of question answering with RAG.
4.2 Implementation Settings
@@
We extract all entities from Freebase, encode their name description into vectors using SBERT model (all-MiniLM-L6-v2, without fine-tuning) (Reimers and Gurevych 2019), and subsequently index them in a vector database (Douze et al. 2024) for downstream entity querying and matching. The knowledge graph derived from Freebase is deployed on a locally deployed Virtuoso (Erling and Mikhailov 2009) server for our experiments. We selected GPT-4o-mini and Qwen-Plus as the fundamental LLM in our framework. To prevent redundant exploration by agents and reduce computational overhead, we set both the maximum search depth $L$ and the expansion width $b$ to 3. This configuration aligns with the empirically optimal values identified in (Sun et al. 2023). Additionally, we fix the number of neighbor relations $k$ during anchor entity grounding to 5, and the number of retrieval agents $m$ is set to 3 for parallel multi-hop reasoning. The data and source can be found in appendix and will be available soon.
4.3 Baselines
@@
Some widely used prompting strategies are selected as the baselines, including vanilla IO prompting (Brown et al. 2020), Chain-of-Thought (CoT) (Wei et al. 2022), and Self-Consistency (Wang et al. 2022). Moreover, we also compare the proposal with two representative RAG methods: Think-on-graph (Sun et al. 2023) and Plan-on-graph (Chen et al. 2024). It is worth noting that the above two methods are both implemented in the closed-world setting where the anchor entities are predefined for each question. Thus, to make a fair comparison, we apply the fundamental LLM to select the anchor entities instead of directly identifying them for the open-world question answering. Other settings and parameters remain unchanged.
4.4 Main Results
The main experimental results, as shown in Table 1, indicate that our proposed method achieves a significant performance advantage over all baseline methods when deployed with both the Qwen-Plus and GPT-4o-mini LLMs. Specifically, when applying Qwen-Plus, AnchorRAG achieves an average improvement by 10% and 11.2% on the metrics Hit@1 and Accuracy respectively, compared to the strongest baseline. Besides, with GPT-4o-mini model, our proposal also achieves the increases of 6.3% and 8.6%. We can also find that our method obtain an outstanding performance on the complex datasets like GrailQA, where its Hit@1 and Accuracy scores are increased substantially by 20.8% and 20.7%, respectively. The reason may be that the questions and the kG entities are semantically inconsistent in this dataset, and the anchor entities are more difficult to locate in this case. AnchorRAG can handle this issue by the multi-agent collaboration, thereby obtaining the state-of-the-art results. In addition, the performance of the RAG methods, e.g., ToG, consistently surpasses the methods solely relying on the internal knowledge, e.g., CoT. This demonstrates the effectiveness of the RAG methods in the question answering task, especially when dealing with the complex reasoning such as GrailQA and CWQ.
@@ 4.4 Main Results
Table 2: Performance comparisons (Hits@1) of different methods on the normal and open-world versions of WebQSP and GrailQA datasets, using Qwen-Plus and GPT-4o-mini as the backbone LLMs.
@@
w/o Entity Grounding 67.8 42.8
w/o Parallel Retrieval 72.8 60.2
w/o Triples Fine Filtering 71.6 62.8
@@
Table 3: Ablation study of AnchorRAG: Effect of different components on the Hits@1 performance.
4.5 Robust Analyses
Given that existing public QA datasets primarily assume that the external KG contains all the necessary information for answering the given questions, which is not suitable for the open-world question answering task. Thus, to further evaluate the generality and robustness of our method, we constructed the open-world versions of the WebQSP and GrailQA datasets. Specifically, we first added the semantic noise by introducing some typos into the original questions to prevent exact matches with entities in the KG. Then we incorporated additional questions into the datasets whose answers could not be directly retrieved from the KG. The detailed construction process is provided in the appendix. The results on the self-constructed datasets are presented in Table 2, we can observe that our method shows strong robustness on this open-world scenario. For instance, on GrailQA, the performance of ToG using Qwen-Plus drops by 23.4%, whereas our method only drops by 11.6%. These results demonstrate that AnchorRAG can effectively address the open-world QA tasks and further enhance the generality of the RAG methods.
4.6 Ablation Studies
@@
To validate the effectiveness of the core components within our proposed AnchorRAG framework, we conducted the following ablation studies on the WebQSP and GrailQA datasets, with the results presented in Table 3. Note that ‘w/o Entity Grounding’ removes the neighborhood information during anchor entity identification, ‘w/o Parallel Retrieval’ eliminates the multi-agent retrieval mechanism, replacing it with a single-path retrieval approach, and ‘w/o Triples Fine Filtering’ refers to the absence of fine semantic filtering for candidate triples during retrieval. The results show that each component can contribute to the question answering performance. Comparing with parallel retrieval and fine filtering, the entity grounding module has a greater impact on the performance, achieving the improvement by 6.3% and 20.6% on WebQSP and GrailQA respectively, which emphasizes the importance of precise anchor entities in the open-world question answering tasks.
Figure 3: The effect of the anchor identifying and knowledge retrieval on WebQSP.
4.7 Effect of the Retrieval

### v2

@@ 4.1. Anchor Entity Identifying
where $\mathcal{N}(e)$ denotes the set of top-$k$ relevant relations according to the Eq.(1). By scoring and ranking all candidates in the set $\tilde{\mathcal{E}}$, the Predictor selects the top-$m$ entities with the higher relevance scores (e.g., France and Nijmegen) as the final anchor set, which will be passed to the following retriever agents for parallel searching.
4.2. Parallel Exploration
With the top-$m$ candidate anchor entities $\{a_{1},a_{2},\dots,a_{m}\}$ obtained by the Predictor agent, the framework initiates a parallel exploration operation, assigning each candidate to an independent retriever agent $\mathsf{R}_{i}$. This multi-agent parallel search is fundamental to the robustness of our framework, as it alleviates the issue of linking to a single, potentially erroneous anchor. Instead of following a linear path, AnchorRAG explores multiple reasoning hypotheses simultaneously. Each retriever agent $\mathsf{R}_{i}$ maintains a reasoning subgraph $G_{i}^{t}=(V_{i}^{t},E_{i}^{t})$ at iteration $t$, initialized as $G_{i}^{(0)}=({a_{i}},\emptyset)$. At each iteration, the agent expands its current exploration frontier by retrieving the one-hop neighborhoods of its active nodes, forming a candidate subgraph. To maintain relevance and reduce overhead, this candidate subgraph then undergoes a two-stage refinement process. First, a candidate ranking step scores and selects the most promising relations and entities for expansion. On this basis, an evidence validation step assesses the generated triples, retaining only those that include relevant evidence for the query. The entities within the validated subgraph are considered as the new frontier nodes for the next iteration of expansion.
Candidate Relations Ranking. At the $t$-th retrieval iteration, each retriever agent dynamically identifies the most relevant relational paths from the active entity $e_{i}^{t}$ through semantic alignment. Given the knowledge graph $\mathcal{G}$, the agent first retrieves all one-hop relations connected to $e_{i}^{t}$, denoted as $\mathcal{R}{(e_{i}^{t})}=\{r_{1},r_{2},\dots,r_{k}\}$.Then a LLM acts as a semantic filter that assigns a relevance score $s_{\text{r}}(r_{j}\mid q,e_{i}^{t})$ to each candidate relation $r_{j}\in\mathcal{R}_{e_{i}^{t}}$. The prompt used for this scoring integrates the query context, the current entity, and all retrieved relations, allowing the model to perform context-aware semantic alignment. Subsequently, the Retriever selects the top-$b$ relations with the highest scores, forming a refined relation set $\tilde{\mathcal{R}}_{e_{i}^{t}}$. As illustrated in Figure 2, for the given query, Retriever $\mathsf{R}_{1}$ starts from the anchor entity $\mathbf{Nijmegen}$” and identifies highly relevant relations such as $\mathbf{nearby\_airports}$”, $\mathbf{serves}$”, and $\mathbf{second\_level\_divisions}$”. Retriever $\mathsf{R}_{2}$ and Retriever $\mathsf{R}_{3}$ start from $\mathbf{France}$”, and then judge that relations like $\mathbf{adjoins}$” and $\mathbf{containedby}$” are beneficial to exploring the answer, while filtering out irrelevant relations such as $\mathbf{contains}$”. During this process, AnchorRAG demonstrates strong generality across diverse knowledge graphs without requiring any training or fine-tuning of models.
Candidate Entities Ranking. At the $t$-th iteration, for each selected relation $r_{j}\in\tilde{\mathcal{R}}{e_{i}^{t}}$, the retriever agent $\mathsf{R}_{i}$ traverses $\mathcal{G}$ to obtain all connected entities, forming a candidate entity set $\mathcal{E}({r_{j}})=\{e_{1},e_{2},\dots,e_{n_{j}}\}$. Note that directly evaluating individual entities without considering their relational context may be suboptimal: such approach may inevitably filter out intermediate entities that are crucial for multi-hop reasoning. To address this problem, $\mathsf{R}_{i}$ employs a LLM-based scoring mechanism that evaluates the contextual relevance $s_{\text{e}}(e_{k}\mid q,e_{i}^{t},r_{j})$ for each entity $e_{k}\in\mathcal{E}_{r_{j}}$. The structured prompt integrates the query $q$, the current anchor entity $e_{i}^{t}$, and the relation $r_{j}$, allowing the model to perform fine-grained semantic reasoning. All candidate entities are then ranked according to their scores, and for each relation $r_{j}$, the top-$b$ entities are retained:
(4) $\tilde{\mathcal{E}}_{r_{j}}=\mathrm{Top}_{b}(\mathcal{E}_{r_{j}},s_{\mathrm{e}})$
These selected entities will form the preliminary reasoning paths $\mathcal{T}(e_{i}^{t})$, which serve as the evidence for the subsequent validation phase.
In addition, to reduce computational cost, we propose an alternative hybrid retrieval strategy that replaces the LLM-based scoring with a combination of sparse (e.g., BM25) and dense (e.g., SBERT) retrieval models. This hybrid approach allows AnchorRAG to efficiently recall semantically or lexically relevant candidates, balancing efficiency and reasoning performance. The effectiveness of this variant is further evaluated in the efficiency evaluation section.
Evidence Validation. In the candidate ranking stage, each candidate relation $r_{j}$ and entity $e_{k}$ is assigned a relevance score $s_{\text{r}}(r_{j})$ and $s_{\text{e}}(e_{k})$, respectively. Existing methods usually apply a linear combination of these two relevance measures to approximate the similarity of the reasoning path formed by the relation–entity pair. They implicitly assume that the relevance of relations and entities is independent, while neglecting their inherent semantic coherence. As a result, they fail to capture the complex semantic interactions and implicit logical structures within the knowledge graph, and thus limit the retrieval performance. To further locate query-relevant facts hidden in the knowledge graph, AnchorRAG introduces a novel evidence validation mechanism. Specifically, each $\mathsf{R}_{i}$ converts its candidate reasoning paths $\mathcal{T}(e_{i}^{t})$ into natural language text sequences, denoted as $\mathcal{X}_{i}^{t}=\{x_{1},x_{2},\dots,x_{p_{i}}\}$, including both the query $q$ and contextual information. On this basis, the agent leverages the reasoning and comprehension capabilities of a LLM to identify relevant paths that contain critical evidence, and filter out irrelevant or inconsistent evidence. As illustrated in Figure 2, Retriever $\mathsf{R}_{1}$ validates the reasoning chain $\mathcal{P}_{1}$ = (Nijmegen, nearby_airports, Weeze Airport), (Weeze Airport, containedby, Germany), while Retriever $\mathsf{R}_{2}$ identifies another path $\mathcal{P}_{2}$ = (France, adjoins, Germany), (Weeze Airport, containedby, Germany). These validated subgraph $\tilde{G}_{i}^{t}$ is then stored in a shared global memory $\mathcal{M}$ accessible to all agents, serving as an important source for subsequent reasoning iterations and answer generation.
@@ 4.2. Parallel Exploration
Table 1. The results of different methods on various Datasets, using Qwen-Plus and GPT-4o-mini as the LLM. The best results are highlighted in bold. The suboptimal results are underlined.
4.3. Reasoning
After the parallel exploration phase, each retriever agent $\mathsf{R}_{i}$ has obtained a reasoning subgraph $G_{i}^{t}$ (and its validated subset $\tilde{G}_{i}^{t}$) which are stored in the shared global memory $\mathcal{M}$. The Supervisor agent $\mathcal{S}$ aggregates these multi-perspective evidences and checks whether the obtained information is sufficient to answer the query $q$. If the evidence chain is complete and results in a clear conclusion, the Supervisor generates the answer directly. If the evidence is insufficient, Supervisor $\mathcal{S}$ evaluates all candidate reasoning paths and drives the high relevance Retriever to explore in a promising direction at the next round, which are marked as $\mathbf{active}]$. Conversely, retrievers deemed to deviate from the query or retrieve noisy, irrelevant information are terminated early, reducing the overhead and limiting the influence of irrelevant data in the global memory. Exploration stops under either of the following two conditions: (1) a predefined maximum search depth $\mathcal{D}$ is reached, or (2) no retrievers remain in the $\mathbf{active}]$ state. In such cases, the Supervisor concludes that further KG-based path exploration is unlikely to generate an answer and applies the chain-of-thought (CoT)(Wei et al., 2022) reasoning mechanism instead to directly obtain the answers.
5. Experiments
@@
Dataset and Evaluation. In the experiments, we adopt four benchmark datasets to evaluate the performance of our framework, including three multi-hop Knowledge Graph Question Answering (QA) datasets: WebQuestionSP (WebQSP) (Yih et al., 2016), GrailQA (Gu et al., 2021) and Complex WebQuestions (CWQ) (Talmor and Berant, 2018), and one for Open-Domain Question Answering: WebQuestions (Berant et al., 2013). Freebase (Bollacker et al., 2008) is utilized as the external Knowledge Graph, which contains the complete knowledge facts that are necessary to answer questions in the above datasets. We select Freebase to ensure fair comparison with established benchmarks(Sun et al., 2023; Chen et al., 2024) whose structural complexity and large scale poses a challenge to evaluating the reasoning capabilities of RAG methods. In addition, we apply the exact match (i.e., Hit@1) and accuracy (i.e., Acc) as evaluation metrics for our method which are widely used in the field of question answering with RAG.
@@
Implementation Settings. We extract all entities from Freebase, encode their name description into vectors using SBERT model (all-MiniLM-L6-v2, without fine-tuning) (Reimers and Gurevych, 2019), and subsequently index them in a vector database (Douze et al., 2024) for downstream entity querying and matching. The knowledge graph derived from Freebase is deployed on a locally deployed Virtuoso (Erling and Mikhailov, 2009) server for our experiments. We selected GPT-4o-mini and Qwen-Plus as the fundamental LLM in our framework. To prevent redundant exploration by agents and reduce computational overhead, we set both the maximum search depth $\mathcal{D}$ and the expansion width $b$ to 3. This configuration aligns with the empirically optimal values identified in (Sun et al., 2023). Additionally, we fix the number of neighbor relations $k$ during anchor entity identifying to 5, and the number of retrieval agents $m$ is set to 3 for parallel multi-hop reasoning.
@@
Baselines. Some widely used prompting strategies are selected as the baselines, including vanilla IO prompting (Brown et al., 2020), chain-of-Thought (CoT) (Wei et al., 2022), and Self-Consistency (Wang et al., 2022). Moreover, we also compare the proposal with two representative RAG methods: Think-on-graph (Sun et al., 2023) and Plan-on-graph (Chen et al., 2024). It is worth noting that the above two methods are both implemented in the closed-world setting where the anchor entities have been pre-defined and are part of the annotated datasets. Thus, to test their performance in the specific scenario where the anchor entity is directly available, we apply the fundamental LLM to select the anchor entities for the following open-world question answering. Other settings and parameters remain unchanged.
@@
Table 2. Performance comparisons (Hits@1) of different methods on the normal and open-world versions of WebQSP and GrailQA datasets, using Qwen-Plus and GPT-4o-mini as the backbone LLMs.
5.1. Main Results
The main experimental results, as shown in Table 1, indicate that our proposed method achieves a significant performance advantage over most baseline methods when deployed with both the Qwen-Plus and GPT-4o-mini LLMs. Specifically, when applying Qwen-Plus, AnchorRAG achieves an average improvement by 10% and 11.2% on the metrics Hit@1 and Accuracy respectively, compared to the strongest baseline. Besides, with GPT-4o-mini model, our proposal also achieves the increases of 6.3% and 8.6%. We can also find that our method obtain an outstanding performance on the complex datasets like GrailQA, where its Hit@1 and Accuracy scores are increased substantially by 20.8% and 20.7%, respectively. The reason may be that the questions and the KG entities are semantically inconsistent in this dataset, and the anchor entities are more difficult to locate in this case. AnchorRAG can handle this issue by the multi-agent collaboration, thereby obtaining the state-of-the-art results. In addition, the performance of the RAG methods, e.g., ToG, consistently surpasses the methods solely relying on the internal knowledge, e.g., CoT. This demonstrates the effectiveness of the RAG methods in the question answering task, especially when dealing with the complex reasoning such as GrailQA and CWQ.
5.2. Robust Analyses
Since existing public QA datasets primarily assume that the external KG contains all the necessary information and rely on clean entity linking, they are ill-suited for the open-world question answering task. Thus, to further evaluate the generality and robustness of our method, we constructed the open-world versions of the WebQSP and GrailQA datasets. Specifically, we first added semantic noise by introducing typos into the original questions. These perturbations can effectively simulate real-world scenarios, thereby testing the capability of the proposed method for the open-world RAG. Then, we incorporated additional questions into the datasets whose answers could not be directly retrieved from the KG. The detailed construction process is provided in the appendix A. The results on the self-constructed datasets are presented in Table 2, where we can observe that our method shows strong robustness on this open-world scenario. For instance, on GrailQA, the performance of ToG using Qwen-Plus drops by 23.4%, whereas our method only drops by 11.6%. These results demonstrate that AnchorRAG can effectively address the open-world QA tasks and further enhance the generality of the RAG methods.
@@ 5.2. Robust Analyses
w/o Entity Disambiguation 67.8 42.8
w/o Parallel Exploration 72.8 60.2
w/o LLM Ranking 71.1 64.2
w/o Evidence Validation 71.6 62.8
@@
Table 3. Ablation study of AnchorRAG: Effect of different components on the Hits@1 performance.
5.3. Ablation Studies
@@
To validate the effectiveness of the core components within our proposed AnchorRAG framework, we conducted the following ablation studies on the WebQSP and GrailQA datasets, with the results presented in Table 3. Note that “w/o Entity Disambiguation” removes the neighborhood information during anchor entity identification, “w/o Parallel Exploration” eliminates the multi-agent retrieval mechanism, replacing it with a single-path retrieval approach. The “w/o LLM Ranking” variant (referred to as AnchorRAG-LR in the efficiency evaluation subsection) uses a lightweight hybrid retrieval strategy instead of LLM ranking. It works by combining SBERT semantic similarity and BM25 scores for efficient candidate filtering. And “w/o Evidence Validation” refers to the absence of fine semantic filtering for candidate triples during retrieval. The results show that each component can contribute to the question answering performance. Comparing with parallel exploration and evidence validation, the entity disambiguation module has a greater impact on the performance, achieving the improvement by 6.3% and 20.6% on WebQSP and GrailQA respectively, which emphasizes the importance of precise anchor entities in the open-world question answering tasks. Interestingly, the “w/o LLM Ranking” variant performs slightly better than the AnchorRAG on GrailQA. This is likely because GrailQA contains more structured and well-aligned relations, where lightweight semantic retrieval (SBERT + BM25) can efficiently the semantic associations without the excessive reasoning conducted by the LLM.
5.4. Efficiency evaluation
In this efficiency evaluation, we compared the performance of AnchorRAG, AnchorRAG-LR along with the strongest baseline ToG in terms of computational efficiency and answer quality. The results are shown in Figure 3, where “Prompt Tokens Consumption (M)” represents the overhead of input tokens that formulate the LLM prompt, “Completion Tokens Consumption (M)” indicates the number of tokens generated in the LLM responses and “Average Inference Time per Sample (s)” measures the average reasoning time for answering each query. From the results, we can obviously observe that AnchorRAG-LR achieves substantially faster reasoning than both ToG and AnchorRAG, while also effectively reducing the overhead of the prompt or completion tokens. For example, on the WebQSP dataset, AnchorRAG-LR reduces the total token consumption by 48.2% compared to ToG and by 77.8% compared to AnchorRAG; its average reasoning time is only 75.3% and 36.2% of that of the methods ToG and AnchorRAG. Similarly, we can observe that on the GrailQA dataset, where token consumption decreases by around 37.4% and 74.2%, and reasoning speed improves by 26.8% and 62.4%, respectively. Additionally, in terms of answer accuracy metrics, AnchorRAG remains superior due to its LLM-based powerful reasoning capability. In summary, AnchorRAG achieves stronger semantic performance at a higher computational cost, whereas AnchorRAG-LR provides a practical and efficient alternative in the resource-constrained scenarios. In addition, we recommend prioritizing the LLM-ranking strategy (AnchorRAG) for those highly ambiguous queries that cannot be easily tackled by lightweight models, e.g., SBERT.
Figure 3. Efficiency and Performance Comparison of Different Methods on the WebQSP and GrailQA Datasets, using GPT-4o-mini as the backbone LLM.
5.5. Effect of the Retrieval